OpenStudy (anonymous):

3x - 4y = 7 3x + 2y = -5 When the second equation is subtracted from the first, the resulting equation is _____. @salehgeek

OpenStudy (anonymous):

\(3x-4y=\) \(7\) \(-(3x+3y=-5)\) ----------------

OpenStudy (anonymous):

its 2y not 3y

OpenStudy (anonymous):

Sorry: \(3x-4y=\) \(7\) \(-(3x+2y=-5)\) ------------------

OpenStudy (anonymous):

Solving this will give you \(-6y=...\)

OpenStudy (anonymous):

-6y= -12 -6y= 12

OpenStudy (anonymous):

Which one did you get?

OpenStudy (anonymous):

-6y= -12

OpenStudy (anonymous):

It should be \(6y=12\) Can you show how you got your answer?

OpenStudy (anonymous):

sorry \(-6=12\)

OpenStudy (anonymous):

3x + 5y = 78 2x - y = 0 The point of intersection of the lines has an x-coordinate of ___.

OpenStudy (anonymous):

sorry \(-6y=12\)

OpenStudy (anonymous):

\(3x-4y=7\) \(-3x-2y=5\) ------------------ \(0-6y=12\)

OpenStudy (anonymous):

x + 5y = -2 2x + y = 5 The point of intersection of the lines has a y-coordinate of ___.

OpenStudy (anonymous):

We rewrite the first equation in terms of x \[x=-2-5y\] Then, we plug the value of x in terms of y into the second formula: \[2(-2-5y)+y=5\] Upon solving for y we have: \[y=-1\]